OpenStudy (anonymous):

Solve, using the substitution method. 4x+2y=11 x–2=–2y

OpenStudy (anonymous):

4(2-2y)+2y=11 8-8y+2y=11 8-6y=11 -6y=3 y=-1/2 put this y in x=2-2y u get, x=2-2(-1/2) x=2+2(1/2) x=2+1 x=3.

OpenStudy (anonymous):

3,-1/2?

OpenStudy (jdoe0001):

one sec

OpenStudy (freckles):

it sounds good @Brandonsnider188 to me :)

OpenStudy (anonymous):

thanks @freckles